OpenStudy (anonymous):

In the past 4 years, a sporting goods store had two yearly losses of $140,000 and $56,000 and two yearly profits of $110,000 and $60,000. What was the net profit or loss over 4 years? A. $196,000 B. $26,000 C. $170,000 D. –$26,000

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

-$140,000 - $56,000 + 110,000 + 60,000 =?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

loses = -140,000-56,000 profit=110,000 +60,000 = -140,000-56,000+110,000 +60,000 =-26,000.

OpenStudy (anonymous):

o i was way off